find the area of the given triangle:
angle A=70 degrees
b=15
c=36

Answer by venugopalramana(3286)

find the area of the given triangle:
angle A=70 degrees
b=15
c=36
AREA =(1/2)bcSIN(A)
=15*36*SIN(70)/2=253.66